What is the mass of an object if it experiences a force of 6.57 N overt a time period of 6.65 while it changes velocity by 11.42m/s
1
Expert's answer
2019-04-05T10:24:50-0400

We can find the mass of the object from the definition of the impulse:


"J = F \Delta t = m \Delta v,"

here, "J" is the impulse, "F" is the force acting on the object, "\Delta t" is the time period during which the force acting on the object, "m" is the mass of the object and "\Delta v" is the change in the velocity.

Then, we get:


"m = \dfrac{F \Delta t}{\Delta v} = \dfrac{6.57 N \cdot 6.65 s}{11.42 \dfrac{m}{s}} = 3.82 kg."

Answer:

"m = 3.82 kg."